Irradiation of chlorine gas (Cl2) to produce chlorine atoms (Cl) in the process described in #1 accelerates the rate at which hydrogen chloride gas is formed. Experimentally, it has been observed that the minimum energy required to produce chlorine atoms in this way is 240 kJ per mole of chlorine. That is, 240 kJ mol-?. In the space provided below, calculate the minimum photon frequency that corresponds to this energy. Report your answer in the box provided with the correct number of significant figures. Show your work.
1
Expert's answer
2020-11-20T10:03:11-0500

Energy required by photons to produce chlorine atoms = 240 KJ/mole

which is equal to (240/NA) KJ/atom where NA is avogadro number.


h"\nu" = "\frac{240000}{N_A}"

"\nu" = "\frac{240000}{N_A×h}"

where NA = 6.023× 1023

and h (plank constant) = 6.626× 10-34 Js


"\nu" = 6.01× 1014 sec-1